• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 23:55:47 on October 15,1998

    Entry number 2916

    Followups:

    keyword=dipoles

    02:00 E_dipole and H_dipole still have not settled to .447 Gev momentum change that was begun during previous shift. Crate rebooted and M.Stevens contacted.